OpenStudy (anonymous):

the "table" shows the change in population for a group of mice over a 4 month period. month: 0 1 2 3 4 pop: 2 8 16 32 64 if this pattern continues, what wll be the population of mice in month 12? a) 192 b) 4096 c) 8192 d) 16384

OpenStudy (amistre64):

2 8 16 32 are the 2^n stuff

OpenStudy (amistre64):

2^0 = 1 so 2^(n+1) looks like a better fit

OpenStudy (anonymous):

what?

OpenStudy (amistre64):

did i miss it again? :)

OpenStudy (anonymous):

no, i just dont get it

OpenStudy (amistre64):

2^1 = 2 2^2 = 4 2^3 = 8 2^4 = 16

OpenStudy (amistre64):

month 0 = 2^(0+1) = 2 month 1 = 2^(1+1) = 4 that one dont fit

OpenStudy (anonymous):

but how does that go with the problem?

OpenStudy (anonymous):

oh, i think i get it now

OpenStudy (amistre64):

if it wasnt for month 1 being 8, itd fit good so I think we might need to ignore that one bad fit

OpenStudy (amistre64):

unless there is a typo in the question :)

OpenStudy (anonymous):

0 doesnt fit either. i dont think that thats it

OpenStudy (anonymous):

no, no typo

OpenStudy (amistre64):

month 2 = 2^(2+r) = 16 = 2^(2+2) = 16 = 2^(n+2) month 12 = 2^(12+2) = 2^14 = 16384 perhaps?

OpenStudy (anonymous):

oh, maybe. ill just ask my teacher tho, thanks!

OpenStudy (amistre64):

k, good luck with it